OpenStudy (anonymous):

At a town fair, 10 tickets cost $8. How many tickets could you expect to receive for $12, assuming there is no discount for buying more tickets?

OpenStudy (anonymous):

@lncognlto

OpenStudy (anonymous):

10/8 = x/12 cross multiply 8x = 120 x = 15

OpenStudy (anonymous):

@anas2000 ty!

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@anas2000 yw

OpenStudy (anonymous):

=)